The ratio of the intensities of two interfering waves is 4 : 1. The ratio of maximum to minimum intensity is

Grade:12th pass

4 Answers

mohammed saud pp
8 years ago
3:1
Adarsh
8 years ago
Given
I1/I2=(a1/a2)2=4/1
a1/a2=2/1
Imax / Imin = (a1+a2)2 / (a1-a2)
               = 9/1
So,  Imax : Imin= 9:1
 
Adarsh
8 years ago
Please approve my answer if it is correct.
 
Yash Chourasiya
3 years ago
Dear Student

Given
I1/I2 = (a1/a2)2 = 4/1
a1/a2 = 2/1
We Know that
Imax/ Imin= (a1 + a2)2/ (a1 - a2)2
= (2 + 1)2/ (2 – 1)2
= 9 / 1
So,Imax: Imin= 9:1

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya